my question is what i did wrong on this problem?
y=x-4
4x+y=26

Answer by stanbon(75887)

y=x-4
4x+y=26
--------------
Substitute y=x-4 into the 2nd equation:
4x+x-4=26
5x=30
x=6
---
Then y=x-4
y=6-2=4
---------
Solution x=6; y=4
----------
Cheers,
Stan H.